# Scalable Multi-Objective and Meta Reinforcement Learning via Gradient Estimation

### v1

@@ Acknowledgments
In the Meta-World experiments, we employ Soft Actor-Critic (SAC) (haarnoja2018soft) as the policy gradient algorithm. SAC is an off-policy, actor-critic deep reinforcement learning algorithm based on the maximum entropy reinforcement learning framework. The core idea is to encourage exploration by maximizing a trade-off between the expected return and the policy’s entropy.
@@
$L(\psi_{i})=\hat{\mathbb{E}}_{(s_{t},a_{t},r_{t},s_{t+1},d_{t})\sim\mathcal{D}}\left[\frac{1}{2}\left(Q_{\psi_{i}}(s_{t},a_{t})-y(r_{t},s_{t+1},d_{t})\right)^{2}\right],$ (28)
@@
$y(r_{t},s_{t+1},d_{t})=r_{t}+\gamma(1-d_{t})\left(\min_{j=1,2}Q_{\psi_{j,\text{target}}}(s_{t+1},a_{t+1})-\alpha\log\pi_{\theta}(a_{t+1}\mid s_{t+1})\right),\text{with }a_{t+1}\sim\pi_{\theta}(\cdot\mid s_{t+1}).$ (29)
The target network parameters $\psi_{j,\text{target}}$ is updated by
@@
In the control environments, we employ Proximal Policy Optimization (PPO) (schulman2017proximal) as our policy gradient algorithm. PPO follows the design of the actor-critic algorithm (konda1999actor), where the core optimization objective is the clipped surrogate policy loss. This is combined with a value function loss and an entropy as the optimization objective.
@@
In the meta-RL setting, we use the MAML algorithm (finn2017model), which consists of a two-step optimization process. In the inner loop adaptation, MAML computes an adapted policy for each source task $T_{i}$ from the trajectory $\mathcal{D}_{\theta}^{i}$ sampled by the interaction with the meta policy $\theta$ and the corresponding environment:
@@
PaCo (Parameter Compositional MTRL) introduces a highly flexible framework where task-specific parameters are explicitly composed rather than separated (sun2022paco). In PaCo, the parameter vector for a given task is constructed as a linear combination of a shared, task-agnostic basis of parameter vectors and a learned, task-specific compositional vector. Other notable architectures, such as Soft Modularization (SM) and those based on Mixture-of-Experts (MoE), explore ways to dynamically route information and learn task-specific components in a unified network (yang2020multi; sun2022paco).
Simultaneous optimization of a shared set of parameters under competing demands of various tasks frequently leads to destructive interference, where joint training results in worse performance on a given task than training it alone (yu2020pcgrad; teh2017distral). The gradients from different tasks can point in conflicting directions, which destabilizes learning (yu2020pcgrad; liu2021cagrad).
@@
Meta reinforcement learning may be viewed as a specific scenario of meta learning. MAML (finn2017model) is a foundational work that introduces a bi-level optimization structure. In the inner loop, the policy is adapted to specific source tasks. In the outer loop, the initial meta-policies are updated by differentiating with respect to inner-loop adaptation, optimizing for the expected performance of the adapted policy across a distribution of source tasks. Building on this, ProMP (rothfusspromp) introduces a low variance curvature estimator to stabilize training by assigning weight to each sample in trajectories. Tr-MAML (collins2020task) focuses on the most challenging task and introduces a min-max framework to minimize the maximum loss in the set.
@@
For the experiment, we maintain a uniform distribution for the target tasks, where goals can appear anywhere on the circle. We then analyze the agent’s adaptation performance after being meta-trained on source tasks drawn from three different distributions. We find that:

### v2

@@ Appendix C Omitted Experimentals
In the Meta-World experiments, we employ Soft Actor-Critic (SAC) [haarnoja2018soft] as the policy gradient algorithm. SAC is an off-policy, actor-critic deep reinforcement learning algorithm based on the maximum entropy reinforcement learning framework. The core idea is to encourage exploration by maximizing a trade-off between the expected return and the policy’s entropy.
@@
$L(\psi_{i})=\hat{\mathbb{E}}_{(s_{t},a_{t},r_{t},s_{t+1},d_{t})\sim\mathcal{D}}\left[\frac{1}{2}\left(Q_{\psi_{i}}(s_{t},a_{t})-y(r_{t},s_{t+1},d_{t})\right)^{2}\right].$ (28)
@@
$y(r_{t},s_{t+1},d_{t})=r_{t}+\gamma(1-d_{t})\left(\min_{j=1,2}Q_{\psi_{j,\text{target}}}(s_{t+1},a_{t+1})-\alpha\log\pi_{\theta}(a_{t+1}\mid s_{t+1})\right),$ (29)
with $a_{t+1}\sim\pi_{\theta}(\cdot\mid s_{t+1})$. The target network parameters $\psi_{j,\text{target}}$ is updated by
@@
In the control environments, we employ Proximal Policy Optimization (PPO) [schulman2017proximal] as our policy gradient algorithm. PPO follows the design of the actor-critic algorithm [konda1999actor], where the core optimization objective is the clipped surrogate policy loss. This is combined with a value function loss and an entropy as the optimization objective.
@@
In the meta-RL setting, we use the MAML algorithm [finn2017model], which consists of a two-step optimization process. In the inner loop adaptation, MAML computes an adapted policy for each source task $T_{i}$ from the trajectory $\mathcal{D}_{\theta}^{i}$ sampled by the interaction with the meta policy $\theta$ and the corresponding environment:
@@
PaCo (Parameter Compositional MTRL) introduces a highly flexible framework where task-specific parameters are explicitly composed rather than separated [sun2022paco]. In PaCo, the parameter vector for a given task is constructed as a linear combination of a shared, task-agnostic basis of parameter vectors and a learned, task-specific compositional vector. Other notable architectures, such as Soft Modularization (SM) and those based on Mixture-of-Experts (MoE), explore ways to dynamically route information and learn task-specific components in a unified network [yang2020multi, sun2022paco].
Simultaneous optimization of a shared set of parameters under competing demands of various tasks frequently leads to destructive interference, where joint training results in worse performance on a given task than training it alone [yu2020pcgrad, teh2017distral]. The gradients from different tasks can point in conflicting directions, which destabilizes learning [yu2020pcgrad, liu2021cagrad].
@@
Meta reinforcement learning may be viewed as a specific scenario of meta learning. MAML [finn2017model] is a foundational work that introduces a bi-level optimization structure. In the inner loop, the policy is adapted to specific source tasks. In the outer loop, the initial meta-policies are updated by differentiating with respect to inner-loop adaptation, optimizing for the expected performance of the adapted policy across a distribution of source tasks. Building on this, ProMP [rothfusspromp] introduces a low variance curvature estimator to stabilize training by assigning weight to each sample in trajectories. Tr-MAML [collins2020task] focuses on the most challenging task and introduces a min-max framework to minimize the maximum loss in the set.
@@
For the experiment, we maintain a uniform distribution for the target tasks, where goals can appear anywhere on the circle. We then analyze the agent’s adaptation performance after being meta-trained on source tasks drawn from three different distributions. The training curves are illustrated in Figure 4(a). We find that meta-training under uniform task distribution consistently outperforms that under partial and biased task distribution. In detail, we find that: